# The Radius Distribution of M dwarf-hosted Planets and its Evolution

### v1

@@ Abstract
M dwarf stars are not only the most promising hosts for detection and characterization of small and potentially habitable planets, they provide leverage relative to solar-type stars to test models of planet formation and evolution. Using Gaia astrometry, adaptive optics imaging, and calibrated gyrochronologic relations to estimate stellar properties, filter binaries, and assign ages, we refined the radii of 179 transiting planets orbiting 119 single late K- and early M-type stars detected by the Kepler mission, and assigned stellar rotation-based ages ) to 115 of these. We constructed the radius distribution planets and assessed its evolution with time. As for solar-type stars, the inferred distribution contains distinct populations of “super-Earths" (at 1.3) and “sub-Neptunes" (at 2.2 ) separated by a gap or “valley" at 1.7that has a period dependence that is significantly weaker (power law index of -0.026) than for solar-type stars. Sub-Neptunes are largely absent at short periods (2 days) and high irradiance, a feature analogous to the “Neptune desert" observed around solar-type stars. The relative number of sub-Neptunes to super-Earths declines between the younger and older halves of the sample (median age 3.8 Gyr), although the formal significance is low () because of the small sample size. The decline in sub-Neptunes appears to be more pronounced at long orbital periods vs. short periods; this is not due to detection bias and could indicate that these objects are inflated by a mechanism that operates at elevated irradiance, e.g. a runaway water greenhouse augmented by H/He.
@@
However, M dwarfs differ from their solar-mass counterparts in ways that could impact the properties and evolution of planets, including an extended pre-main sequence phase, longer rotational spin-down times, and elevated X-ray and extreme ultraviolet emission relative to bolometric emission. There may also be differences in their protoplanetary disks, i.e. their masses (Gaidos, 2017), lifetimes (Silverberg et al., 2020), and chemistry (Tabone et al., 2023). To assess potential variation in planetary evolution, precise measurement of the properties of a large number of planets in systems with robust ages that span a wide age range is needed. Currently, the confirmed and candidate transiting planets among the Kepler Objects of Interest (KOIs) provide the largest, most uniform, and most precisely measured set of planet properties, i.e., radii and periods (e.g., Thompson et al., 2018).
The radius distribution of sub-Jovian planets around solar-type stars detected by the Kepler mission contains two distinct peaks, representing populations of “super-Earths"- and “sub-Neptunes" separated by a gap or “valley" at 1.7 (Owen & Wu, 2013; Fulton et al., 2017; Berger et al., 2020). Complementary measurements of planet mass (and hence density) indicate that sub-Neptunes and super-Earth have rocky cores with and without envelopes of volatile, low molecular-weight constituents (Weiss & Marcy, 2014; Rogers, 2015). These envelopes could consist of hydrogen and helium captured from the primordial gas disk during planet formation (e.g., Béthune & Rafikov, 2019), and/or H2O and other volatiles acquired as ices (Turbet et al., 2020; Luque & Pallé, 2022). Unfortunately, mass measurements of small planets are observationally expensive, often not always precise, and the different mass-radius relations are at least partly degenerate (e.g., Rogers & Seager, 2010).
@@
Reconstructing the planet radius distribution of M dwarf planets and its evolution requires radii and ages that are more precise and accurate than have previously been available. Transiting planet radius depends on host star radius, and calculations of the radius distribution that account for detection efficiency require stellar density (both mass and radius). New empirical calibrations for M dwarfs permit estimation of mass and radius based on the precise absolute luminosities made possible by Gaia parallaxes (Mann et al., 2015, 2019). Moreover, both Gaia astrometry and ground-based adaptive optics (AO) allow more thorough vetting for stellar companions, which can affect the radius estimate. Binary stars also appear to have a planet population distinct from single stars (Kraus et al., 2016; Sullivan et al., 2023).
For field M dwarfs, rotation (gyrochronology) is the most promising method to assign ages compared to other techniques (i.e., isochrone-fitting, asteroseismology, depletion of lithium, or kinematics), but in the absence of accurate, first-principles spin-down models, this approach requires empirical calibration. Dungee et al. (2022) constructed a rotation sequence (vs. effective temperature ) for late K-type and early-type M dwarfs in the 4 Gyr-old M67 cluster, allowing calibrated age-dating of 3200–4200K stars to nearly half the age of the Galactic disk. Moreover, a comparison with younger clusters suggests these stars are spinning down according to a simple power-law prescription (Skumanich, 1972) by ages of a few Gyr, justifying an extrapolation to older ages. Gaidos et al. (2023b) constructed -dependent age-rotation relations using rotation sequences and assigned ages to the cool host stars of known exoplanets, including Keplerconfirmed and candidate planets.
In this work we revise the radius distribution of Kepler planets around single late K- and early M-type dwarf stars and examine its variation with age, taking advantage of (1) precise Gaia parallaxes for luminosity calculations and revised mass/radius/luminosity relations based on nearby stars; (2) identification and screening of binary systems using astrometry and AO imaging; and (3) assignment of robust rotation-based ages based on a new gyrochronology (Gaidos et al., 2023a). We compare the M dwarf planet radius distribution to that derived for solar-type stars and interpret it and its change with time to different scenarios for planet composition and evolution.
@@ 2 Analysis
We matched the Kepler Input Catalog (KIC, Brown et al., 2011) with Gaia Data Release 3 (DR3) sources (Gaia Collaboration et al., 2023) using an angular separation criterion of 0".71. Absolute -band (2.2 µm) magnitudes () were computed using photometry from the 2MASS Point Source Catalog (Skrutskie et al., 2006; Cutri et al., 2013) and parallaxes from Gaia Data Release 3 (Gaia Collaboration et al., 2016, 2023). We restricted our analysis to stars observed by Kepler during Quarters 1-17 with (inclusive of spectral types later than K5, Pecaut & Mamajek, 2013), and that appear in the catalog of late K and M-type dwarfs for which Santos et al. (2019) obtained rotation signals from Kepler lightcurves (about 59% of the 24,162 stars they analyzed). We further limited the target sample to stars with values in Santos et al. (2019) between 3200 and 4200K, the limit of the gyrochronology which we apply to these stars (Gaidos et al., 2023b), yielding a final sample of 5433 target stars that Kepler observed with the relevant host star parameters. Around these stars we identified objects from the DR25 version of the Kepler Objects of Interest (KOI) catalog based on data from Quarters 1-17 (Thompson et al., 2018), excluding KOIs with Exoplanet Archive designations of “false positive".111To maximize uniformity and the robustness of the statistics, we did not incorporate the DR25 Supplemental Catalog, which includes a few planets (candidates) erroneously categorized as false positives in DR25. https://exoplanetarchive.ipac.caltech.edu/docs/PurposeOfKOITable.html#q1-q17_sup_dr25 Finally, we restricted the sample to KOIs with gyrochronologic age assignments from Gaidos et al. (2023b).
Host star masses were calculated from an empirical mass-luminosity relationship in (Eqn. 4 in Mann et al., 2019). Radii were calculated from using Eqns. 4 or 5 in Mann et al. (2015) for stars with or without estimates of [Fe/H] retrieved from the Exoplanet Archive.222These [Fe/H] estimates are typically from published analyses, to be distinguished from the metallicity estimates included in the original KIC. Errors were estimated using Monte Carlo calculations assuming normally-distributed errors in -band magnitudes and parallaxes.
@@ 2.1 Filtering of binary stars
Stars in binary or multiple systems with separations 50 au are known to host a planet population that differs from that around single stars (Kraus et al., 2016). Furthermore, binary stars with separations (100 au) have different rotational histories than their single counterparts (e.g., Messina et al., 2017; Stauffer et al., 2018), meaning that a rotation-age relation developed for single stars cannot be applied. 90% of M dwarfs in the Kepler field are more distant than 135 pc, thus these physical separations correspond to angular separations 1". Other effects of companions or any nearby sources unresolved or partially resolved by Kepler (which had 4" pixels and a 3.1-7.5" PSF) are that the transit signals are diluted and identification of the host star itself – and thus planet properties – may be ambiguous. The -band photometry used to estimate radius and mass will be inaccurate for binaries that are unresolved by the 2MASS survey (resolution of 3"). Thus we identified and excluded KOIs around binary stars with separations ". Binary/multiple systems or candidate systems were identified and excluded using (1) Gaia astrometry of resolved sources; (2) Gaia astrometric error relative to single-star astrometric solutions; (3) adaptive optics imaging; and (4) over-luminosity relative to a single-star locus in a color-absolute magnitude diagram. Here the procedures and outcomes are summarized; details will be published in a forthcoming paper.
Gaia resolved companions: We identified companions resolved by Gaia (1 arcsec, Ziegler et al., 2018a) based on similarity of parallax (distance) and proper motion. Rather than fixed criteria for similarity (e.g., El-Badry et al., 2021), we adopted a Bayesian approach. We identified candidate companions with cone searches of the DR3 catalog with a radius of 74" (corresponding to about au at the median distance of Kepler M dwarfs) around each Kepler M dwarf. We computed the probability of companionship as , where and are the likelihoods that the source is a companion or background star, respectively. We evaluated these assuming normally-distributed errors using:
@@
and are the absolute scalar difference in proper motion between the star and source and its standard error, and are the difference in parallax and its standard error, and is a “softening" term that accounts for the angular orbital motion of a binary. This last term we set to 5 mas yr-1 based on Eqn. 4 in El-Badry & Rix (2018). Constant factors that divide out are omitted from Eqns. 1 and 2. The difference in the terms involving parallax and proper motion reflects the latter’s two-dimensionality. To improve the statistical accuracy by averaging over many stars, the summation for is performed over the stars in all the other cone searches (a few thousand stars) that cannot be companions. We imposed no weighting with angular separation other than the generous search radius. Based on the observed distribution of values relative to those calculated using stars from other fields, we identified a threshold of to select companions.
Gaia astrometric error: Binaries with separations " cannot be resolved by Gaia (Ziegler et al., 2018b) but can often be identified based on the poor fit of Gaia astrometry to a single-star solution (Belokurov et al., 2020). This is quantified as the Reduced Unit Weight Error (RUWE, Lindegren et al., 2018) and values significantly greater than unity are indicative of a multiple star system. We selected a threshold of 1.4 based on the distribution RUWE values of AO-identified binaries (Kraus et al., in prep.).
AO imaging: AO imaging of KOIs was obtained of 75 of 86 systems (116 of 139 planets) with the NIRC2 infrared imager and AO system on the Keck-2 telescope (McLean & Chaffee, 2000). Observations used the narrow camera (10 mas pixel scale) and either the or filters. Images were processed, sources extracted, and astrometry performed using the pipeline described in Kraus et al. (2016). We selected those sources with separations 2" and contrast as highly likely to be bound companions based on Kraus et al. (2016).
@@ 2.2 Planet radius distribution
The posterior distribution in radius of each planet was determined by multiplying each value from the Monte Carlo Markov Chain (MCMC) posterior chains for planet-to-star radius ratio from DR25 Q1-17 (Hoffman & Rowe, 2017) with a value from the stellar posterior distribution in radius calculated above, assuming Gaussian errors in magnitude and parallax. We normalized each distribution by the summed probability for detection of an equivalent planet around every star observed by Kepler that is a late K- or early M-type dwarf , 3200 4200K) and that appears in the Santos et al. (2019) catalog of stars evaluated for rotation periods. We used the KeplerPORTs code (Burke & Catanzarite, 2017b) to calculate the probability of detection (including the geometric transit probability). The required window functions, one-sigma detection functions (Burke & Catanzarite, 2017a), “dataspans" (timespans of data), and duty cycles were retrieved from the NASA Exoplanet Archive. Non-linear (four-coefficient) limb-darkening coefficients were obtained from the DR25 Stellar Properties Catalog (Mathur et al., 2017). To mimic the removal of binaries from the KOI list, we excluded all target stars with Gaia RUWE values 1.4. This distilled the catalog to 182 KOIs in 122 systems. Finally, KOIs were then matched with entries in the TIME Table of Gaidos et al. (2023b) to retrieve gyrochronologic age estimates. We identified 115 KOIs (76 systems) with entries in the tabulation of Gaidos et al. (2023a) with a median age of 3.86 Gyr. The summed probabilities over all stars represent the expected number of detections if each star in the observed sample had such a planet. Each posterior distribution in radius was divided by this quantity and then summed to give the number of objects per unit radius per star. As a metric of uncertainty due to radius error and finite counting statistics, we calculated 68% confidence intervals by constructing 1000 Monte Carlo realizations using bootstrap sampling of the planet population with replacement.444While multi-planet systems will affect the counting statistics of total planet occurrence (which is not considered here), correlation between the radii of planets in the same system (e.g., Weiss et al., 2018) would impact the significance of individual features.

### v2

@@ Abstract
M dwarf stars are the most promising hosts for detection and characterization of small and potentially habitable planets, and provide leverage relative to solar-type stars to test models of planet formation and evolution. Using Gaia astrometry, adaptive optics imaging, and calibrated gyrochronologic relations to estimate stellar properties and filter binaries we refined the radii of 117 Kepler Objects of Interest (confirmed or candidate planets) transiting 74 single late K- and early M-type stars, and assigned stellar rotation-based ages to 113 of these. We constructed the radius distribution of 115 small (4) planets and assessed its evolution. As for solar-type stars, the inferred distribution contains distinct populations of “super-Earths" (at 1.3) and “sub-Neptunes" (at 2.2 ) separated by a gap or “valley" at 1.7that has a period dependence that is significantly weaker (power law index of -0.03) than for solar-type stars. Sub-Neptunes are largely absent at short periods (2 days) and high irradiance, a feature analogous to the “Neptune desert" observed around solar-type stars. The relative number of sub-Neptunes to super-Earths declines between the younger and older halves of the sample (median age 3.86 Gyr), although the formal significance is low () because of the small sample size. The decline in sub-Neptunes appears to be more pronounced on wider orbits and low stellar irradiance. This is not due to detection bias and suggests a role for H2O as steam in inflating the radii of sub-Neptunes and/or regulating the escape of H/He from them.
@@
However, M dwarfs differ from their solar-mass counterparts in ways that could impact the properties and evolution of their planets, including an extended pre-main sequence phase (Luger & Barnes, 2015), longer rotational spin-down times (e.g., Johnstone et al., 2021), and elevated X-ray and extreme ultraviolet emission relative to bolometric emission (e.g., France et al., 2016). There may also be differences in their protoplanetary disks, i.e. their masses (Gaidos, 2017), lifetimes (Silverberg et al., 2020), and chemistry (Tabone et al., 2023). To assess potential variation in planetary evolution, precise measurement of the properties of a large number of planets in systems with robust ages that span a wide age range is needed. Currently, the confirmed and candidate transiting planets among the Kepler Objects of Interest (KOIs) provide the largest, most uniform, and most precisely measured set of planet properties such as radii and periods (e.g., Thompson et al., 2018).
The radius distribution of sub-Jovian planets around solar-type stars detected by the Kepler mission contains two distinct peaks, representing populations of “super-Earths"- and “sub-Neptunes" separated by a gap or “valley" at 1.7 (Owen & Wu, 2013; Fulton et al., 2017; Berger et al., 2020). Complementary measurements of planet mass (and hence density) indicate that sub-Neptunes and super-Earth have rocky cores with and without envelopes of volatile, low molecular-weight constituents (Weiss & Marcy, 2014; Rogers, 2015). These envelopes could consist of hydrogen and helium captured from the primordial gas disk during planet formation (e.g., Béthune & Rafikov, 2019), and/or H2O and other volatiles acquired as ices (Turbet et al., 2020; Luque & Pallé, 2022). Unfortunately, mass measurements of small planets are observationally expensive, often not precise, and the mass-radius relations are at least partly degenerate with composition (e.g., Rogers & Seager, 2010).
@@
Reconstructing the planet radius distribution of M dwarf planets and its evolution requires radii and ages that are more precise and accurate than have previously been available. Transiting planet radius depends on host star radius, and calculations of the radius distribution that account for detection efficiency require stellar density (both mass and radius). New empirical calibrations for M dwarfs permit estimation of mass and radius based on the precise absolute luminosities made possible by Gaia parallaxes (Mann et al., 2015, 2019). Moreover, both Gaia astrometry and ground-based adaptive optics (AO) imaging allow more thorough vetting for stellar companions, which, if not properly accounted for, can affect a radius estimate. Binary stars also appear to have a planet population distinct from single stars (Kraus et al., 2016; Sullivan et al., 2023).
For field M dwarfs, rotation (gyrochronology) is the most promising method to assign ages compared to other techniques (i.e., isochrone-fitting, asteroseismology, depletion of lithium, or kinematics), but in the absence of accurate, first-principles spin-down models, this approach requires empirical calibration. Dungee et al. (2022) constructed a rotation sequence (vs. effective temperature ) for late K-type and early-type M dwarfs in the 4 Gyr-old M67 cluster, extended calibrated age-dating of 3200–4200K stars to nearly half the age of the Galactic disk. Moreover, a comparison with younger clusters suggests these stars are spinning down according to a simple power-law prescription (Skumanich, 1972) by ages of a few Gyr, justifying an extrapolation to older ages. Gaidos et al. (2023) constructed -dependent age-rotation relations using rotation sequences and assigned ages to the cool host stars of known exoplanets, including Keplerconfirmed and candidate planets.
In this work we revise the radius distribution of Kepler Objects of Interest (KOIs, i.e., confirmed, validated or candidate planets) around single late K- and early M-type dwarf stars and examine its variation with age, taking advantage of precise Gaia parallaxes for luminosity calculations and revised mass/radius/luminosity relations based on nearby stars; identification and screening of binary systems using astrometry and AO imaging; and assignment of robust rotation-based ages based on the Gaidos et al. (2023) gyrochronology. We compare the M dwarf planet radius distribution to that derived for solar-type stars and interpret it and its change with time to different scenarios for planet composition and evolution.
@@ 2 Analysis
We matched the Kepler Input Catalog (KIC, Brown et al., 2011) with Gaia Data Release 3 (DR3) sources (Gaia Collaboration et al., 2023) using an angular separation criterion of 0".71. Absolute -band (2.2 µm) magnitudes () were computed using photometry from the 2MASS Point Source Catalog (Skrutskie et al., 2006; Cutri et al., 2013) and parallaxes from Gaia Data Release 3 (Gaia Collaboration et al., 2016, 2023). We restricted our analysis to stars observed by Kepler during Quarters 1-17 with (inclusive of spectral types later than K5, Pecaut & Mamajek, 2013), and that appear in the Santos et al. (2019) catalog of late K and M-type dwarfs with rotation signals from Kepler lightcurves (about 59% of the 24,162 stars analyzed). We further limited the sample to stars with values in Santos et al. (2019) between 3200 and 4200K, the valid limits of the Gaidos et al. (2023) gyrochronology to be applied to these stars , yielding a final sample of 5433 target stars that Kepler observed with the relevant host star parameters. We identified KOIs around these stars from the DR25 catalog, which is based on data from Quarters 1–17 (Thompson et al., 2018).
We excluded KOIs that have false positive (FP) dispositions in the DR25 Supplemental Catalog, which includes a few planets (candidates) erroneously categorized as false positives in DR25.111https://exoplanetarchive.ipac.caltech.edu/docs/PurposeOfKOITable.html#q1-q17_sup_dr25. We also excluded those with a disposition of “candidate" and a modified astrophysical FP probability calculated with vespa (Morton, 2015). This threshold was chosen based on the distributions of FP probability; above , FPs outnumber confirmed or validated planets. The modified probability is the sum of the probabilities of the undiluted eclipsing binary (EB) and hierarchical EB scenarios, plus the background EB scenario if our AO imaging cannot exclude this. The latter entails that imaging was obtained, that there are no additional sources detected within 2" of the KOI (and thus falling on the same Kepler pixel and inducing minimal centroid motion, Adams et al., 2012) and that the contrast limit of the imaging is sufficient to detect an EB sufficiently bright to produce the observed signal. For this we adopt a maximum eclipse depth of 50% (zero impact parameter) and a Kepler- reddening of 0.24 for background sources. The latter is based on a mean total reddening for the Kepler field (Schlafly & Finkbeiner, 2011) and the extinction coefficients of Yuan et al. (2013), substituting Sloan for the Kepler bandpass. Hereafter we refer to all KOIs that pass these tests as “planets".
Host star masses were calculated from an empirical mass-luminosity relationship (Eqn. 4 in Mann et al., 2019). Radii were calculated from using Eqns. 4 or 5 in Mann et al. (2015) for stars with or without estimates of [Fe/H] retrieved from the Exoplanet Archive.222These [Fe/H] estimates are typically from published analyses, to be distinguished from the metallicity estimates included in the original KIC. Errors were estimated using Monte Carlo calculations assuming normally-distributed errors in -band magnitudes and parallaxes. Finally, we identified those KOIs with gyrochronologic age assignments in Gaidos et al. (2023).
@@ 2.1 Filtering of binary stars
Stars in binary or multiple systems with separations 50 au are known to host a planet population different from that around single stars (Kraus et al., 2016). Furthermore, binary stars with separations (100 au) have different rotational histories than their single counterparts (e.g., Messina et al., 2017; Stauffer et al., 2018), meaning that a rotation-age relation developed for single stars cannot be applied. 90% of M dwarfs in the Kepler field are more distant than 135 pc, thus these physical separations correspond to angular separations 1". Companions or any nearby sources unresolved or partially resolved by Kepler (which had 4" pixels and a 3.1-7.5" PSF) can also dilute transit signals and make identification of the host star itself – and thus planet properties – ambiguous. The -band photometry used to estimate radius and mass will be inaccurate for binaries that are unresolved by the 2MASS survey (resolution of 3"). Thus we identified and excluded KOIs around binary stars with separations ". Binary/multiple systems or candidate systems were identified and excluded using (1) Gaia astrometry of resolved sources; (2) Gaia astrometric error relative to single-star astrometric solutions; (3) adaptive optics imaging; and (4) over-luminosity relative to a single-star locus in a color-absolute magnitude diagram.
Gaia resolved companions: We identified companions resolved by the Gaia satellite (1 arcsec, Ziegler et al., 2018a) based on similarity of parallax (distance) and proper motion. Rather than fixed criteria for similarity (e.g., El-Badry et al., 2021), we adopted a Bayesian approach. We identified candidate companions with cone searches of the DR3 catalog with a radius of 74" (corresponding to about au at the median distance of Kepler M dwarfs) around each Kepler M dwarf. We computed the probability of companionship as , where and are the likelihoods that the source is a companion or background star, respectively. We evaluated these assuming normally-distributed errors using:
@@
where and are the absolute scalar difference in proper motion between the star and source and its standard error, and are the difference in parallax and its standard error, and is a “softening" term that accounts for the angular orbital motion of a binary. This last term we set to 5 mas yr-1 based on Eqn. 4 in El-Badry & Rix (2018). Constant factors that divide out are omitted from Eqns. 1 and 2. The difference in the terms involving parallax and proper motion reflects the latter’s two-dimensionality. To improve the statistical accuracy by averaging over many stars, the summation for is performed over the stars in all the other cone searches (a few thousand stars) that cannot be companions. We imposed no weighting with angular separation other than the generous search radius. Based on the observed distribution of values relative to those calculated using stars from other fields, we identified a threshold of to select companions.
Gaia astrometric error: Binaries with separations " cannot be resolved by Gaia (Ziegler et al., 2018b) but can often be identified based on the poor fit of Gaia astrometry to a single-star solution (Belokurov et al., 2020). This is quantified as the Reduced Unit Weight Error (RUWE, Lindegren et al., 2018) and values significantly greater than unity are indicative of a multiple star system. We selected a threshold of 1.4 based on Lindegren et al. (2021).
AO imaging: AO imaging of KOIs was obtained of all but 8 systems with the NIRC2 infrared imager and AO system on the Keck-2 telescope (McLean & Chaffee, 2000). Observations used the narrow camera (10 mas pixel scale) and either the or filters. Images were processed, sources extracted, and astrometry performed using the pipeline described in Kraus et al. (2016). We selected those sources with separations 2" and contrast as highly likely to be bound companions based on Kraus et al. (2016).
@@ 2.2 Planet radius distribution
The posterior distribution in radius of each planet was determined by multiplying each value from the Monte Carlo Markov Chain (MCMC) posterior chains for planet-to-star radius ratio from DR25 Q1-17 (Hoffman & Rowe, 2017) with a value from the stellar posterior distribution in radius calculated above, assuming Gaussian errors in magnitude and parallax. To correct for detection efficiency, we normalized each distribution by the summed probability for detection of an equivalent planet around every star observed by Kepler that is a late K- or early M-type dwarf , 3200 4200K) and that appears in the Santos et al. (2019) catalog of stars evaluated for rotation periods. We used the KeplerPORTs code (Burke & Catanzarite, 2017b) to calculate the probability of detection (including the geometric transit probability). The required window functions, one-sigma detection functions (Burke & Catanzarite, 2017a), “dataspans" (timespans of data), and duty cycles were retrieved from the NASA Exoplanet Archive. Non-linear (four-coefficient) limb-darkening coefficients were obtained from the DR25 Stellar Properties Catalog (Mathur et al., 2017). To mimic the removal of binaries from the KOI list, we excluded all target stars with Gaia RUWE values 1.4. Finally, KOIs were then matched with entries in the TIME Table of Gaidos et al. (2023) to retrieve gyrochronologic age estimates.
The summed probabilities over all stars represent the expected number of detections if each star in the observed sample had such a planet. Each posterior distribution in radius was divided by this quantity and then summed to give the number of objects per unit radius per star. As a metric of uncertainty due to radius error and finite counting statistics, we calculated 68% confidence intervals by constructing 1000 Monte Carlo realizations using bootstrap sampling of the planet population with replacement.444While multi-planet systems will affect the counting statistics of total planet occurrence (which is not considered here), correlation between the radii of planets in the same system (e.g., Weiss et al., 2018) would impact the significance of individual features.